OpenStudy (anonymous):

Can someone help me with this question pls and thanks in advance. The equation y=-1777x+27153 can be used to predict the number y of gun deaths in the united states x years after 2000, that is x=0 corresponds to 2000,x=0 corresponds to 2002 x=3 corresponds to 2003 x=5 and so on. predict the number of gun deaths in 2005 and 2007. In what year will the number of gun deaths be 11,160?

OpenStudy (anonymous):

your posting this question over and over again

OpenStudy (anonymous):

This is my first time on here and the first time, I mixed the numbers up. I thought I deleted it. Sorry

OpenStudy (anonymous):

okay for the first part, if 2002 x=3 and 2003, x=5, then this means that 2005, x=9 and 2007, x=13. So for 2005 simply evaluate this :y=-1777(9)+27153 For 2007 evaluate this:y=-1777(13)+27153 and for the last part: solve this: 11160=-1777x+27153

OpenStudy (anonymous):

Thank you very much